OpenStudy (anonymous):

how much money should be invested at 5% compounded monthly to produce $10,000 in 7years?

OpenStudy (jdoe0001):

\(\bf Balance= P\left(1+\frac{r}{n}\right)^{nt}\\ \quad \\ r = \textit{rate in decimal format, }5\% = 0.05\\ P= \textit{Principal or original amount}\\ n= \textit{periods per year, 12 in 1year}\\ t = years\) solve for "P" and plug in your values

OpenStudy (jdoe0001):

notice that \(\bf 10,000= P\left(1+\frac{0.05}{12}\right)^{12\cdot 7}\)

OpenStudy (anonymous):

do i divide 10,000 from the other side?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

yes, that'd be one way, yes

OpenStudy (wolf1728):

5 % compounded monthly equals (1.05/12)^12 =1.0511618979 = 5.11618979% yearly interest We need a formula that solves for principal Principal = Total ÷ (1 + Rate)^Years Principal = 10,000 ÷ (1.0511618979)^7 Principal = 10,000 ÷ 1.4180360522 Principal = 7,052.01